P(A)=0.68P(A)=0.68, P(B)=0.27P(B)=0.27 and P(AandB)=0.05P(AandB)=0.05.

P(notB∣notA)=
Answer #1

P(A) = 0.68
P(B) = 0.27
P(A and B) = 0.05
P(not A) = 1 - P(A) = 1 - 0.68 = 0.32
P(not A and not B) = 1 - P(A or B) = 1 - [P(A) + P(B) - P(A and B)] = 1 - (0.68 + 0.27 - 0.05) = 0.1

Therefore using bayes theorem now we get:

P(not B | not A) = P(not A and not B) / P( not A) = 0.1 / 0.32 = 0.3125

Therefore 0.3125 is the required probability here.